Assume that adults have IQ scores that are normally distributed with a mean of 99.2 and a standard deviation of 24.8. Find the probability that a randomly selected adult has an IQ greater than 125.8.
Answer #1

Mu=99.2, sigma=24.8

P(X > 125.8) = P(Z>z)

Z = (X-mu)/sigma = ( 125.8-99.2)/24.8 =1.07

P ( Z > 1.07 ) = 1 - P ( Z < 1.07 ) = 1- 0.85769 =0.14231